# Revisiting Cross-Diffusion for Overcrowding Dispersal in Interacting Species System

### v1

@@ Abstract
This work introduces and study a novel class of reaction-diffusion systems to model the evolution of two interacting species that disperse to avoid overcrowding. Our approach employs the principle of proximal minimization energy, implemented through a minimum flow proximal algorithm. This framework offers a potential generalization of existing systems commonly used in the theory of segregation for interacting species dynamics. Beyond existence and uniqueness, this approach allows us to capture the dynamic interplay between diffusion rates and concentration gradients of each species, and has the potential to significantly advance our understanding of how cross-diffusion shapes the spatial distribution, coexistence, and pattern formation of multiple species within a system.
@@ 1 Introduction and preliminaries
Fo reach $k=1,2,$ let us denote by $\rho_{k}(t,x),$ the concentration of each specie $k,$ at time $t\geq 0$ at position $x\in\Omega$, where $\Omega\subset I\!\!R^{N}$ is a given open bounded domain. Here $N\geq 1$ may be considered equals to $1,2$ or 33 33 for practical situations. A particular example of cross-diffusion system we shall discuss in this paper for the couple of interacting densities $(\rho_{1},\rho_{2})$ is given by the balances law for each $\rho_{k},$ $k=1,2,$
$\partial_{t}\rho_{k}+\nabla\cdot J_{k}[\rho_{1},\rho_{2}]=f_{k}$
where $J_{k}[\rho_{1},\rho_{2}]$ is the mass-flux and $f_{k}$ represents reaction term or simply species supply. To model the dispersal phenomena Gurtin et al. propose in [15] to consider the intuitive constitutive relation for each $J_{k}$ given by
@@
where $0<\sigma_{k}=\sigma_{k}(\rho_{k},\rho_{1}+\rho_{2})$ is assumed to be connected to the dispersivity, so that each specie disperses (locally) toward lower values of total population. Then, the underlying PDE takes the form of cross-diffusion system
@@
(1.1) $\left\{\begin{array}[]{ll}\partial_{t}\rho_{1}-\nabla\cdot(\rho_{1}\>\sigma_{1}(\rho_{k},\rho_{1}+\rho_{2})\>\nabla(\rho_{1}+\rho_{2}))=f_{1}\\ \\ \partial_{t}\rho_{2}-\nabla\cdot(\rho_{2}\>\sigma_{2}(\rho_{k},\rho_{1}+\rho_{2})\>\nabla(\rho_{1}+\rho_{2}))=f_{2}.\end{array}\right.$
@@
Since the seminal paper by [15], the cross-dffusion system (1.1) was proposed in a series of works with constant dispersivity parameter assumed to be equal to 11 11 ; i.e;
@@
(1.2) $\left\{\begin{array}[]{ll}\partial_{t}\rho_{1}-\nabla\cdot(\rho_{1}\>\nabla(\rho_{1}+\rho_{2}))=f_{1}\\ \\ \partial_{t}\rho_{2}-\nabla\cdot(\rho_{2}\>\nabla(\rho_{1}+\rho_{2}))=f_{2}.\end{array}\right.$
to study segregation theory (see [7, 8, 10, 16]).
See that, while the system initially appears to exhibit fast-wave propagation (hyperbolic behavior) for each individual density $\rho_{k},$ it surprisingly becomes diffusion-dominated (parabolic behavior) when considering the total cell density $\rho_{1}+\rho_{2}$ in the specific case where both mobilities $\sigma_{1}$ and $\sigma_{2}$ are equal to a single value dependent on the total density $\sigma(\rho_{1}+\rho_{2})$. Interestingly, the segregation property suggests that the system might still retain some diffusion-like behavior even for individual densities.
Our study explores through a new approach how two interacting species, driven to disperse by overcrowding, influence each other’s spatial distribution. This investigation reveals a previously hidden diffusion-like (parabolic) nature of the underlying dynamics, offering valuable new insights into their behavior. Notably, this allows us to address well-posedness questions, such as existence and uniqueness of solutions for more general situations. Moreover, while the hyperbolic formulation presented in equation (1.1) may appear suitable, it is not ideal for scenarios involving the dynamic with Dirichlet boundary conditions, such as crowd motion at an exit. This is because the hyperbolic nature suggests rapid wave propagation, which might not accurately capture the dynamics of controlled outflow through a specific location like an exit. As we will explore further, this presents an intriguing challenge that our approach seeks to address.
We introduce a parabolic version of the dynamic, contrasting with the traditional approaches linked to [15] which models mass fluxes using phenomenological processes. The key difference lies in how we handle the mass flux, $J_{k}[\rho_{1},\rho_{2}]$ for each species $k=1,2.$ Our approach, however, utilizes an optimization framework based on the system’s internal energy, expressed through the total density. Inspired by work on single-species scenarios [19], we employ a steepest descent process to minimize this energy against the diffusion work of each species. This framework allows us to introduce the concept of dispersal potentials, $\eta_{k}.$ These potentials capture the individual’s desire to migrate based on the local densities of both species and the total density, $\rho_{1}+\rho_{2}.$ In other words, we encode the dispersal behavior using specific parameters, $\eta_{k}=\eta_{k}(\rho_{1},\rho_{2},\rho_{1}+\rho_{2}),$ to rewrite the mass flux for each species directly as:
$J_{k}[\rho_{1},\rho_{2}]=-\nabla\eta_{k},\quad\hbox{ in }Q,\hbox{ for each }k=1,2.$
@@
$\left\{\begin{array}[]{l}\partial_{t}\rho_{k}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k})=f_{k},\\ \\ \eta_{k}=\eta_{k}(\rho_{1},\rho_{2},\rho_{1}+\rho_{2})\end{array}\right.\quad\text{in }Q,\text{ for }k=1,2.$
@@
By delving deeper into the optimization process for the system’s internal energy, we uncover a crucial link between the dispersal potentials, $\eta_{k},$ the individual densities, $\rho_{k},$ and the total density, $\rho_{1}+\rho_{2}.$ This connection leads into significant equations, providing valuable tools for analyzing the dynamic behavior of the system. Beyond the standard state equation for nonlinear diffusion phenomena (observed in porous media, Stefan, and Hele-Shaw problems), a unique relationship emerges definitely between the total density and the maximum of the dispersal potentials, $\max(\eta_{1},\eta_{2}).$ Additionally, the dispersal potentials satisfy specific equations for spatial delimitation of region:
(1.3) $\rho_{1}=0\text{ in }[\eta_{1}<\eta_{2}]\quad\text{and}\quad\rho_{2}=0\text{ in }[\eta_{1}>\eta_{2}].$
In particular, combining these equations with the state equation enable the description of dynamics within the drivers regions (see Remark 2:
  • $[\rho_{1}>0\>\&\>\rho_{2}=0]$ and $[\rho_{1}=0\>\&\>\rho_{2}>0]$ : regions exclusively populated by $\rho_{1}$ and $\rho_{2}$, respectively.
  • $[\rho_{1}>0\>\&\>\rho_{2}>0]$: region potentially occupied by both species.
It’s important to note that the regions $[\eta_{1}>\eta_{2}]$ and $[\eta_{2}>\eta_{1}]$ represents the so-called segregated or separated regions, while $[\eta_{1}=\eta_{2}]$ corresponds integrated or desegregated region. This novel approach offers a unique perspective on dispersal dynamics, leveraging an optimization framework to capture the interplay between species and their surrounding environment. It provides a powerful tool for analyzing and understanding the emergence of spatial segregation and the dynamics within these distinct regions.
@@
This framework leverages an internal energy optimization approach instead of direct mass flux modeling. This enables us to effectively connect dispersal potentials, individual species densities, and the total density within the dynamic system. It underpins further analysis in this study. Notably, it enables us to investigate the general scenario where each species dynamics are governed together by their intrinsic overcrowding dispersal forces, external forces acting like transport/convection phenomena and various boundary conditions, including Dirichlet, Neumann, or mixed, and not necessary homogeneous, for each population. To this aim, we introduce the cross-diffusion system
@@
where $\eta_{k}=\eta_{k}(\rho_{1},\rho_{2},\rho_{1}+\rho_{2}),$ may be interpreted as dispersal potentials, satisfy conditions (1.3) and need to fill moreover the state equations
(1.5) $\max(\eta_{1},\eta_{2})\in\partial\beta(\rho_{1}+\rho_{2}),$
where $\partial\beta$ represents the sub-differential operator of $\beta,$ which is a given intrinsic application closely connected to the internal energy (see Section 2). Moreover, they enables to handle different boundary conditions for each specie like
(1.6) $\left\{\begin{array}[]{ll}(\sigma_{k}\>\nabla\eta_{k}+\rho_{k}V_{k})\cdot\nu=\pi_{k}&\hbox{ on }\Sigma_{N_{k}}:=(0,T)\times\Gamma_{N_{k}}\\ \eta_{k}=g_{k},&\hbox{ on }\Sigma_{D_{k}}:=(0,T)\times\Gamma_{N_{k}}\end{array}\right.$
where $\Gamma_{N_{k}}$ and $\Gamma_{N_{k}}$ constitute partitions of the boundary $\partial\Omega$, and the couples $(g_{k},\pi_{k})$ are given boundary data.
The plan of the paper is the following: in Section 2, we gather some notation and the assumptions and state the main results. We present the existence of a weak solution for the problem (3.37). Then, for the homogeneous case and vanishing transport; i.e. $V\equiv 0,$ we show how one can connect the problem (3.37) to the well established $H^{-1}-$gradient flow dynamic. In Section 3, we show how to establish the model by using minimum flow steepest gradient descent algorithm. In Section 4, we prove the existence of a weak solution by employing steepest gradient descent algorithm and passing to the limit in the approximate solution. Finally, in Section 5, we provide the proof of the connection of the approach with $H^{-1}-$gradient flow in in the homogeneous case.
@@ 2 Assumptions and main results
$\partial\Omega=\underbrace{\Gamma_{D_{1}}\cup\Gamma_{D_{2}}}_{\Gamma_{D}}\cup\underbrace{\Gamma_{N_{1}}\cup\Gamma_{N_{2}}}_{\Gamma_{N}},$
such that
@@
Moreover, we fix $\alpha=(\alpha_{1},\alpha_{2})\in I\!\!R^{2}$ and $\sigma=(\sigma_{1},\sigma_{2})\in[L^{\infty}_{+}(\Omega)]^{2}$ such that
(2.7) $0<\min(\alpha_{1},\alpha_{2})\quad\hbox{ and }\quad 0<\min_{x\in\overline{\Omega}}\min(\sigma_{1}(x),\sigma_{2}(x)).$
Our aim is to introduce and study the cross-diffusion system
(2.8) $\left\{\begin{array}[]{ll}\partial_{t}\rho_{k}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k}+\rho_{k}V_{k})=f_{k},\quad\hbox{ for }k=1,2\\ \\ \rho_{1},\>\rho_{2}\geq 0,\>\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})\\ \\ \rho_{1}\left(\frac{\eta_{2}}{\alpha_{2}}-\frac{\eta_{1}}{\alpha_{1}}\right)^{+}=\rho_{2}\left(\frac{\eta_{1}}{\alpha_{1}}-\frac{\eta_{2}}{\alpha_{2}}\right)^{+}=0\end{array}\right\}\quad\hbox{ in }Q,$
Here, $\vee$ denotes the maximum operator. $V_{k}$ represents a vector-valued field, and $f_{k}$ is a source term, assumed to be provided for both species. We consider system (2.8) with mixed boundary conditions of the following type:
@@
where $g_{k}$ and $\tilde{\pi}_{k}$ are given boundary data, for $k=1,2,$ we precise later.
@@
    $\|z\|_{H^{1}(\Omega)}=\left(\int_{\Omega}|z|^{2}+\int_{\Omega}|\nabla z|^{2}\right)^{1/p}.$
    $H^{1}_{\tilde{\Gamma}}(\Omega)$ denotes the closure, in $H^{1}(\Omega)$, of $\mathcal{C}^{1}(\overline{\Omega})$ which are null on $\tilde{\Gamma}.$ In particular, for any $z\in H^{1}_{\tilde{\Gamma}}(\Omega),$ $\gamma(z)=0$ on $\tilde{\Gamma},$ where $\gamma$ denotes the usual trace application defined from $H^{1}(\Omega)\to L^{2}(\partial\Omega).$
    $H^{-1}_{\tilde{\Gamma}}(\Omega)$ denotes the dual space of $H^{1}_{\tilde{\Gamma}}(\Omega).$ We know that (see for instance [19]) , for any $f\in H^{-1}_{\tilde{\Gamma}}(\Omega),$ there exists a couple $(f_{0},\overline{f})\in L^{2}(\Omega)\times L^{2}(\Omega)^{N},$ such that
@@
    $H^{1/2}(\tilde{\Gamma})$ denotes the usual dual Sobolev space given by $\gamma_{/\tilde{\Gamma}}(H^{1}(\Omega)),$ where $\gamma_{/\tilde{\Gamma}}$ is the trace application restricted to $\tilde{\Gamma}.$ We need to define moreover
@@
    Thanks to the assumptions on $\partial\Omega,$ taking $\tilde{\Gamma}=\Gamma_{D_{k}}$ or $\tilde{\Gamma}=\Gamma_{N_{k}},$ for $k=1,2,$ the space $H^{1/2}_{00}(\tilde{\Gamma})$ coincides with the set of functions belonging to $H^{1/2}(\tilde{\Gamma})$ and vanishing on the boundary $\partial\Omega\setminus\tilde{\Gamma}.$ Remember that this not automatically true for any $\tilde{\Gamma}\subset\partial\Omega$.
    $H^{-1/2}(\tilde{\Gamma})$ denotes the usual the topological dual space of $H^{1/2}_{00}(\tilde{\Gamma}),$ $\left(H^{1/2}_{00}(\tilde{\Gamma})\right)^{\prime}.$ We denote the duality bracket simply by the formal expression $\langle\pi,\kappa\rangle_{\tilde{\Gamma}}$ ; i.e.

### v2

# Cross-Diffusion Theory for Overcrowding Dispersal in Interacting Species System

@@ Abstract
This work introduces a new class of cross-diffusion systems for studying overcrowding dispersal of two species. The approach, based on proximal minimization energy through a minimum flow process, offers a potential generalization of existing segregation models. Unlike prior methods using PDEs or $W_{2}$-Wasserstein flows, it establishes a well-posed PDE framework for capturing the interplay between diffusion and concentration gradients. This framework has the potential to significantly improve our understanding of how cross-diffusion shapes spatial patterns, coexistence, and overall distribution of multiple species. Notably, for homogeneous cases, the approach definitely leads to a well-defined PDE grounded in a new general $H^{-1}$-theory specifically developed for overcrowding dispersal. This theory provides a robust foundation for further analysis.
@@ 1 Introduction and preliminaries
Let us denote by $\rho_{k}(t,x)$ the concentration of each specie $k=1,2,$ l at time $t\geq 0$ at position $x\in\Omega$, where $\Omega\subset I\!\!R^{N}$ is a given open bounded domain. Here $N\geq 1$, which represents space dimension, may be considered equals to $1,2$ or 33 33 for practical situations. A particular example of cross-diffusion system we shall discuss in this paper for the couple of interacting densities $(\rho_{1},\rho_{2})$ is given by the balances law for each $\rho_{k},$ $k=1,2,$
$\partial_{t}\rho_{k}+\nabla\cdot J_{k}[\rho_{1},\rho_{2}]=f_{k},\quad\hbox{ in }Q:(0,T)\times\Omega,$
where $T>0$ is a given horizon time, $J_{k}[\rho_{1},\rho_{2}]$ is the mass-flux and $f_{k}$ represents reaction term or simply species supply. To model the dispersal phenomena Gurtin et al. propose in [15] to consider the intuitive constitutive relation for each $J_{k}$ given by
@@
where $0<\sigma_{k}$ is assumed to be connected to the dispersivity, and may depend on $\rho_{k}$ and $\rho_{1}+\rho_{2},$ so that each specie disperses (locally) toward lower values of total population. Then, the underlying PDEs take the form of cross-diffusion system
@@
(1.1) $\left\{\begin{array}[]{ll}\partial_{t}\rho_{1}-\nabla\cdot(\sigma_{1}\>\rho_{1}\>\nabla(\rho_{1}+\rho_{2}))=f_{1}\\ \\ \partial_{t}\rho_{2}-\nabla\cdot(\sigma_{2}\>\rho_{2}\>\nabla(\rho_{1}+\rho_{2}))=f_{2}\end{array}\right.\quad\hbox{ in }Q.$
@@
Since the seminal paper by [15], the cross-dffusion system (1.1) was proposed in a series of works with $\sigma_{k}=1$ ; i.e
@@
(1.2) $\left\{\begin{array}[]{ll}\partial_{t}\rho_{1}-\nabla\cdot(\rho_{1}\>\nabla(\rho_{1}+\rho_{2}))=f_{1}\\ \\ \partial_{t}\rho_{2}-\nabla\cdot(\rho_{2}\>\nabla(\rho_{1}+\rho_{2}))=f_{2}.\end{array}\right.\quad\hbox{ in }Q$
subject to homogeneous Neumann boundary condition to study segregation theory where one species dominates (see [8, 9, 11, 16]). Following the works in [7, 11, 21] in the framework of gradient flow in Wasserstein space it is well-established that systems like (1.1) are closely linked to a minimization process of an internal energy functional acting on the total mass $\rho_{2}+\rho_{2},$ i.e. $\int\tilde{\beta}(\rho_{1}+\rho_{2})$. However, existing results are often partial and heavily influenced by ”stable” configurations (as defined in [21]) that avoid mixing between the species. This limitation restricts the study of these type of systems to specific spatial dimensions and source/reaction term behaviors $f_{k}.$ The case where $f_{k}$ arises from a drift is concerned as well, since the study become particularly complex even in one dimensional, as explored in [21].
Actually, while the system (1.1) initially appears to exhibit fast-wave propagation (hyperbolic behavior) for each individual density $\rho_{k},$ it surprisingly becomes diffusion-dominated (parabolic behavior) when considering the total cell density $\rho_{1}+\rho_{2}$. Interestingly, stable configuration (the segregation regime ; i.e. $\rho_{1}\>\rho_{2}=0$) suggests that the system might still retain some diffusion-like behavior even for individual densities. Nevertheless, this strong formulation (hyperbolic) in (1.2) essentially implies that the dynamics of each population are solely determined by the combined dynamics of both populations. This seems to imposes a somewhat inflexible constraint on the individual dynamic of each population. The situation appears to be better from theoretical point of view when the reaction terms $f_{k}$ make possible segregation phenomena (see [8, 9, 11, 16]). The formulation is suitable also when coming moreover with a diffusion terms for each populations allowing to consider in some sens individual Brownian motion for each populations (cf. [22] for the study of crowd motion). In this case the mass-fluxes turns into
$J_{k}[\rho_{1},\rho_{2}]=-\sigma_{k}\>\nabla(\rho_{1}+\rho_{2})-\nabla\rho_{k},$
so that a diffusion terms $-\Delta\rho_{k}$ associated with each density $\rho_{k}$ enter in the system and make it more flexible from the theoretical and numerical points of view.
Our approach in this paper is different, we follow gradient flow approach in Wasserstein spaces, by using the system’s internal energy expressed through a cost functional acting on the total density ; $\int_{\Omega}\beta(\rho_{1}+\rho_{2}).$ Here $\beta$ is assumed to be a convex and lower semi-continuous function, which may vary spatially. We employ a process based on the steepest descent algorithm. However, to measure the transition work, we utilize the minimal flow approach we introduced in [18] for single-species, as an alternative of Wasserstein distance. Roughly speaking, for a given initial condition $(\rho_{01},\rho_{02})$ and a fixed time step $\tau>0,$ we set $(\rho^{1}_{0},\rho_{2}^{0})=(\rho_{01},\rho_{02}),$ and then recursively define
(1.3) $(\rho_{1}^{i},\rho_{2}^{i}):=\hbox{argmin}_{(\rho_{1},\rho_{2})}\left\{\frac{1}{2\tau}(\mathcal{I}(\rho_{1},\rho_{1}^{i-1})+\mathcal{I}(\rho_{2},\rho_{2}^{i-1})+\int_{\Omega}\beta(\rho_{1}+\rho_{2})\>dx\right\},\quad i=1,...n,$
where $\mathcal{I}(\rho_{k},\rho_{k}^{i-1})$ is given by the minimum flow problem associated with the densities $\rho_{k}$ and $\rho_{k}^{i-1}$ (the exact definition in Section 3.3). Remember that this minimal flow approach is well-suited for incorporating diffusion processes into dynamical systems which naturally tend towards minimizing their internal energy (cf. [18]). This framework allows us here to recover the concept of dispersal potentials, $\eta_{k},$ which capture the individual’s desire to migrate based on the local densities of both species and the total density $\rho_{1}+\rho_{2}.$ Basically, we encode the dispersal behavior using specific parameters, $\eta_{k},$ to rewrite the mass flux for each species directly as:
$J_{k}[\rho_{1},\rho_{2}]=-\sigma_{k}\>\nabla\eta_{k},\quad\hbox{for each }k=1,2.$
@@
$\partial_{t}\rho_{k}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k})=f_{k},\quad\text{ for }k=1,2.$
@@
By delving deeper into the optimization process for the system’s internal energy, we uncover a crucial link between the dispersal potentials, $\eta_{k},$ the individual densities, $\rho_{k},$ the total density, $\rho_{1}+\rho_{2},$ and the spatial rearrangements of the densities. This connection leads into significant equations, providing valuable tools for analyzing the dynamic behavior of the system. For instance, in the case where $\beta$ is differentiable, this link may be written in the following form
(1.4) $\eta_{k}-\beta^{\prime}(\rho_{1}+\rho_{2})\in\partial I\!\!I_{[0,\infty)}(\rho_{k})\quad\hbox{ for each }k=1,2,$
which is equivalent to say (see Remark 1 for different related equivalent formulations)
(1.5) $\left\{\begin{array}[]{ll}\eta_{k}=\beta^{\prime}(\rho_{1}+\rho_{2})&\hbox{ in}[\rho_{k}>0]\\ \\ \eta_{k}\leq\beta^{\prime}(\rho_{1}+\rho_{2})&\hbox{ in }[\rho_{k}=0].\end{array}\right.$
In the context of the dispersal overcrowding example (1.2), this novel approach introduces a new model to capture the dynamics of this phenomenon. The method utilizes a parabolic function linked to a minimum internal energy process acting on the total density, which in this case is squared; specifically, this integral is given by : $\frac{1}{2}\int_{\Omega}(\rho_{1}+\rho_{2})^{2}.$ The approach differs from traditional methods rooted in [15] which rely on phenomenological processes to model mass fluxes. The key distinction lies in how we handle the mass flux, $J_{k}[\rho_{1},\rho_{2}],$ for each species $(k=1,2).$ We propose a different treatment within their active zones and outside them. While the proposed system enables to cover the system (1.2) in the segregated regions (i.e. $[\rho_{1}>0\ \&\>\rho_{2}=0]$ and $[\rho_{1}=0\ \&\>\rho_{2}>0]$), it fundamentally changes in the regions where neither species are present (i.e., $[\rho_{1}>0\ \&\>\rho_{2}>0]$). More precisely, we propose reformulating the fluxes in (1.2) as follows
(1.6) $\left.\begin{array}[]{ll}J_{k}[\rho_{1},\rho_{2}]=-(\rho_{1}+\rho_{2})\>\nabla(\rho_{1}+\rho_{2})-\nabla\tilde{\eta}_{k}\\ \\ \tilde{\eta}_{k}\in\partial I\!\!I_{[0,\infty)}(\rho_{k})\end{array}\right\}\quad\hbox{ for }k=1,2.$
Here $\tilde{\eta}_{k}$ is an unknown potential to be determined also. From mathematical point of view it can be viewed as a Lagrange multiplier associated with the positive sign constraint of the density $\rho_{k}.$ From a physical perspective, it could be interpreted as a corrective potential which prevents the system from collapsing into an unphysical two-phase problem (i.e., changing sign solutions). This offers valuable significant insights into the behavior of the system. Notably, this allows us to address well-posedness questions, such as existence and uniqueness of solutions for more general situations (see Section 4).
Additionally, although the hyperbolic formulation in equation (1.2) might seem appropriate for Neumann boundary conditions (conservative case), it becomes less suitable for scenarios involving dynamic behavior with Dirichlet boundary conditions, such as collective motions with distinct ”evacuation exits”. This is because the hyperbolic nature suggests rapid wave propagation, which might not accurately capture the dynamics of controlled outflow through a specific location like an exit. As we will explore further, among others this presents an intriguing challenge that our approach seeks to address. Actually, following the approach [18], by incorporating the boundary condition in an adequate way in the internal energy (see Section 3.3), we suggest to handle the associate boundary value problem through the general mixed boundary conditions
$\left\{\begin{array}[]{ll}J_{k}[\rho_{1},\rho_{2}]\cdot\nu=\pi_{k}&\hbox{ on }\Gamma_{N_{k}}\\ \\ \eta_{k}=g_{k}&\hbox{ on }\Gamma_{N_{k}}\end{array}\right.\quad\text{ for }k=1,2,$
where, for each $k=1,2,$ $\Gamma_{N_{k}}$ and $\Gamma_{D_{k}}$ constitute partitions of the boundary $\partial\Omega,$ $\pi_{k}$ represents a given normal flux for each specie $k$ and $g_{k}$ is a given charge related to Dirichlet boundary section (an exit for instance).
@@
Despite the variety of boundary conditions, the framework enables also to integrate further analysis in the study of overcrowding dispersal dynamics. Notably, it allows to investigate the general scenario where each specie dynamic is governed together by their intrinsic overcrowding dispersal forces and external actions acting like transport/convection phenomena. To this aim, we introduce the cross-diffusion system
@@
where we require that the dispersal potentials, $\eta_{k},$ depend on the local densities of both species ($\rho_{1}$ and $\rho_{2}$), as well as the total density $\rho_{1}+\rho_{2},$ following the specified treatment (1.4), or equivalently, any of the formulations presented in Remark 1. Here $V_{k}$ is a given drift for each specie ($k=1,2$) assumed to be in $L^{\infty}(Q).$
Beside the freedom the new system demonstrates for well-posedness questions for describing overcrowding dispersal dynamics, the approach establishes that in the homogeneous case ; i.e. $\pi_{k}=0$ and $g_{k}=0$, the underlying PDE falls within a new general $H^{-1}$-gradient flow theory specifically developed for overcrowding dispersal. In fact, the transition works $\mathcal{I}(\rho_{k},\rho_{k}^{i-1})$ in (1.3) we consider coincides with the square of $H^{-1}$-norm of $\rho_{k}-\rho_{k}^{i-1}$ in this specific case (see Section 5). This theoretical framework provides a robust foundation for further analysis of the system’s behavior like uniqueness and large time behavior.
It is meaningful to mention that while gradient flow theories in $H^{-1}$ and $W_{2}$-Wasserstein space share many conceptual similarities for single-species overcrowding, existing studies for multi-species system using $W_{2}$-Wasserstein for interacting species haven’t yielded a valuable PDE framework for well-posedness. Additionally, for specific scenarios like segregated densities, the $W_{2}$-Wasserstein approach seems to align more closely with hyperbolic à la Gurtin et al. models than the parabolic cross-diffusion system proposed here. This suggests a need for further investigation into the connections between these approaches for cross-diffusion systems.
The most related work appears to be by [21], who explore degenerate cross-diffusion models with different drift velocities within a $W_{2}$-Wasserstein framework. In this work, the authors consider the case where the energy is a power $m$ of the total density $\rho_{1}+\rho_{2}$ with $m$ possibly equal to infinity. While their approach shares some similarities with ours at the discrete-time level, the complexity of handling density mixing and strong restrictions both on the drift and the space dimension limit the applicability of their continuum limit. This highlights the potential of our proposed approach for broader applicability in studying overcrowding dispersal dynamics.
The plan of the paper is the following: in Section 2, we gather some notation and the assumptions and state the main results. We present the existence of a weak solution for the problem (3.51). Then, for the homogeneous case and vanishing transport; i.e. $V\equiv 0,$ we show how one can connect the problem (3.51) to a well established $H^{-1}-$gradient flow dynamic. In Section 3, we show how to establish the model by using minimum flow steepest gradient descent algorithm. In Section 4, we prove the existence of a weak solution by employing steepest gradient descent algorithm and passing to the limit in the approximate solution. Finally, in Section 5, we provide the proof of the connection of the approach with $H^{-1}-$gradient flow in the homogeneous case.
@@ 2 Assumptions and main results
(2.8) $\partial\Omega=\underbrace{\Gamma_{D_{1}}\cup\Gamma_{D_{2}}}_{\Gamma_{D}}\cup\underbrace{\Gamma_{N_{1}}\cup\Gamma_{N_{2}}}_{\Gamma_{N}},$
with
@@
Moreover, we fix $\sigma=(\sigma_{1},\sigma_{2})\in[L^{\infty}(\Omega)]^{2}$ such that
(2.9) $0<\min_{x\in\overline{\Omega}}\min(\sigma_{1}(x),\sigma_{2}(x)).$
In line with the preliminary discussion in the introduction, we will now formally introduce and analyze the cross-diffusion system :
(2.10) $\left\{\begin{array}[]{ll}\partial_{t}\rho_{k}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k}+\rho_{k}V_{k})=f_{k},\\ \\ \eta_{k}-\tilde{\eta}\in\partial I\!\!I_{[0,\infty)}(\rho_{k}),\>\tilde{\eta}\in\partial\beta(x,\rho_{1}+\rho_{2})\end{array}\right\}\quad\hbox{ in }Q,\quad\hbox{ for }k=1,2.$
Here, $V_{k}$ represents a vector-valued field, and $f_{k}$ is a source term, assumed to be provided for both species. We consider system (2.10) with mixed boundary conditions of the following type:
@@
where $g_{k}$ and $\tilde{\pi}_{k}$ are given boundary data, for $k=1,2,$ we precise later. We will provide a detailed motivation for this system in Section 3.
@@
    Moreover, $L^{p}_{+}(\Omega)$ denotes the set of nonegative functions of $L^{p}(\Omega).$
@@
    $\|z\|_{H^{1}(\Omega)}=\left(\int_{\Omega}|z|^{2}+\int_{\Omega}|\nabla z|^{2}\right)^{1/2}.$
    $H^{1}_{\tilde{\Gamma}}(\Omega)$ denotes the closure, in $H^{1}(\Omega)$, of $\mathcal{C}^{1}(\overline{\Omega})$ functions which are null on $\tilde{\Gamma}.$ In particular, for any $z\in H^{1}_{\tilde{\Gamma}}(\Omega),$ $\gamma(z)=0$ on $\tilde{\Gamma},$ where $\gamma$ denotes the usual trace application defined from $H^{1}(\Omega)\to L^{2}(\partial\Omega).$
    $H^{-1}_{\tilde{\Gamma}}(\Omega)$ denotes the dual space of $H^{1}_{\tilde{\Gamma}}(\Omega),$ $(H^{1}_{\tilde{\Gamma}}(\Omega)^{\prime}.$ We know that (see for instance [18]) , for any $f\in H^{-1}_{\tilde{\Gamma}}(\Omega),$ there exists a couple $(f_{0},\overline{f})\in L^{2}(\Omega)\times L^{2}(\Omega)^{N},$ such that
@@
    Notice that the couple $(f_{0},\overline{f})$ is not unique. However, the resulting dual bracket does not depend on this specific choice.
    $H^{1/2}(\tilde{\Gamma})$ denotes the usual space given by $\gamma_{/\tilde{\Gamma}}(H^{1}(\Omega)),$ where $\gamma_{/\tilde{\Gamma}}$ is the trace application restricted to $\tilde{\Gamma}.$ We need to define moreover
@@
    That is $f=f_{0}+\nabla\cdot\overline{f}$ in $\mathcal{D}^{\prime}(\overline{\Omega}\setminus\tilde{\Gamma}).$ Thanks to (2.8), taking $\tilde{\Gamma}=\Gamma_{D_{k}}$ or $\tilde{\Gamma}=\Gamma_{N_{k}},$ for $k=1,2,$ the space $H^{1/2}_{00}(\tilde{\Gamma})$ coincides with the set of functions belonging to $H^{1/2}(\tilde{\Gamma})$ and vanishing on the boundary $\partial\Omega\setminus\tilde{\Gamma}.$ Remember that this not automatically true for any $\tilde{\Gamma}\subset\partial\Omega$.
    $H^{-1/2}(\tilde{\Gamma})$ denotes the usual (topological) dual space of $H^{1/2}_{00}(\tilde{\Gamma}),$ $\left(H^{1/2}_{00}(\tilde{\Gamma})\right)^{\prime}.$ We denote the duality bracket simply by the formal expression $\langle\pi,\kappa\rangle_{\tilde{\Gamma}}$ ; i.e.